OpenStudy (anonymous):

For each positive integer n, prove that (1+1/n)^{n}<(1+1/n+1)^{n}

OpenStudy (jamesj):

What you've written isn't true. But what is true is that for all positive integers n, \[ \left(1 + \frac{1}{n}\right)^n < \left(1 + \frac{1}{n+1}\right)^{n+1} \]